# From Flatland to Space: Teaching Vision-Language Models to Perceive and Reason in 3D

### v5

@@ 4.1 Evaluation on existing 2D spatial benchmarks
CV-Bench [34].
The Cambrian Vision-Centric Benchmark (CV-Bench) is a 2,638-sample dataset for evaluating 2D spatial reasoning and object counting. It repurposes vision datasets with natural language questions and uses visual prompts (e.g., bounding boxes) to reduce ambiguity and ensure reliable evaluation.
VSI-Bench [20].
The Visual-Spatial Intelligence Benchmark (VSI-Bench) evaluates spatial reasoning from egocentric video data, containing 5,000+ QA pairs from 288 real-world videos sourced from ScanNet [28], ScanNet++ [29], and ARKitScenes [35]. It covers diverse indoor environments and leverages object-level annotations for question generation. VSI-Bench includes eight tasks across three categories: configurational reasoning (object counting, relative position, route planning), measurement estimation (object and room size, absolute distance), and spatiotemporal reasoning (appearance order in videos). All questions undergo manual review to ensure accuracy.
To evaluate the effectiveness of our dataset, we compare models trained on SPAR-mix (ours with EMOVA-2M) against those trained solely on EMOVA-2M. Since our dataset focuses specifically on spatial perception and reasoning, we incorporate a subset of QA data from EMOVA to maintain performance on general-purpose vision-language tasks. Note that the Base VLM has already been pretrained on extensive data with optimized training strategies, making it an unfair baseline for assessing the incremental gain. Therefore, we consider Base VLM + EMOVA-2M as the proper baseline. These results validate the effectiveness of our dataset design and indicate that spatial supervision can be introduced without compromising general QA performance. As shown in table 2, our model achieves state-of-the-art results on CV-Bench 3D and VSI-Bench, demonstrating significant improvements in spatial understanding. While there are minor drops on general benchmarks such as MMBench and TextVQA—mostly within 1%—this is expected, as our data emphasizes spatial reasoning rather than fine-grained vision-language alignment.
VSI-Bench [20] CV-Bench   [34] BLINK [18] 3DSRBench [19] Seed-Image [21] MME [36] MMBench [24] RealWorldQA [37] TextVQA [38]
@@
GPT-4v [39] - 64.3 73.8 51.14 - 71.6 1927 75.0 61.4 77.4
GPT-4o [39] 34.0 - - 60.04 45.3 77.1 2310 83.4 75.4 -
Cambrain-8B [34] - 72.3 72.0 - - 74.7 1547 75.9 64.2 77.8
LLaVA-OV-7B [4] 25.3 - - 48.2 44.1 - 1998 80.8 66.3 -
InternVL2-8B [40] 34.6 - - - - 75.4 2215 81.7 - 77.4
@@
Table 2: Comparison of model performance on various spatial understanding and 2D general benchmarks. The base VLM is InternVL2.5-8B, and we use dynamic patch of 3. Note: Our main comparison is between +EMOVA-2M and +SPAR-mix, as both build upon the same Base VLM and general QA data. For detailed discussion, see section 4.1.
QA source Hypersim [41] SUNRGBD [42] nuScenes [43]
@@
We also find that pretraining with our dataset enhances the model’s spatial understanding across diverse, non-homogeneous datasets. Although our dataset is primarily derived from indoor scene datasets such as ScanNet [28], ScanNet++ [29], and Structured3D [30], the pretrained model shows consistent improvements across all tasks constructed with datasets different from ours in CV-Bench 3D, as shown in table 3.
Notably, our method achieves an 18-point gain on the outdoor nuScenes dataset, highlighting strong generalization beyond indoor scenes. This improvement stems from the spatial reasoning skills emphasized in our dataset—many CV-Bench 3D questions focus on relative distance comparisons, similar to tasks like DistI-OO and ObjRel. Since these tasks rely on 3D geometric cues rather than appearance or domain-specific semantics, the learned spatial priors transfer effectively across diverse environments.
@@ Methods
GPT-4o [39] 3 36.39 29.25 53.80 45.00 15.00 13.60 37.40 34.40 23.40 24.40 24.93 30 16 28.80 45.11 64 64 58 46 46 32 44 30 22
Claude-3.7-Sonnet [44] 5 21.77 25.43 41.00 45.40 11.20 12.20 42.60 19.60 26.00 5.40 7.33 16 6 0.00 23.33 40 48 22 36 14 12 20 6 12
Qwen2-VL-72B [45] 4 35.62 35.28 45.40 49.80 13.80 10.00 54.60 49.40 36.80 22.40 23.39 42 18 10.16 40.00 60 68 50 38 44 18 28 18 36
Qwen2.5-VL-72B [46] 2 39.40 35.35 53.20 46.80 17.80 29.00 49.60 57.40 14.40 14.60 23.05 40 16 13.16 48.44 74 74 60 56 50 20 34 24 44
@@
InternVL2-2B [40] 10 28.06 21.74 18.06 24.81 23.20 20.97 19.47 19.95 26.83 20.61 22.83 39.69 23.00 5.81 35.42 51.18 55.95 46.00 31.59 23.82 36.02 34.30 17.55 22.41
InternVL2-4B [40] 4 32.01 28.94 23.94 27.22 20.00 18.12 42.57 40.16 31.29 28.18 29.16 49.87 21.00 16.62 35.70 56.76 55.36 40.25 36.81 25.21 28.76 32.27 21.19 24.65
InternVL2-8B [40] 3 33.02 26.83 25.75 30.88 20.67 20.78 39.03 36.19 19.15 22.19 36.49 63.36 28.00 18.11 37.37 64.71 54.46 42.75 37.36 26.32 34.14 31.10 20.86 24.65
InternVL2.5-2B [47] 8 30.14 25.79 39.67 39.72 12.12 15.03 30.94 29.59 20.22 19.02 22.93 37.91 24.25 6.64 36.41 51.47 56.85 50.25 33.79 24.10 27.15 35.17 26.49 22.41
InternVL2.5-4B [47] 7 30.55 25.66 29.06 32.97 21.77 16.83 20.84 26.85 28.13 28.79 29.75 47.07 33.25 8.92 35.16 54.12 58.93 35.50 29.67 34.63 24.73 31.39 19.21 28.29
InternVL2.5-8B [47] 1 36.28 29.46 25.78 29.31 23.79 18.76 46.82 42.68 22.62 25.89 31.88 61.32 28.00 6.32 43.80 59.71 56.85 51.75 44.23 41.55 36.56 41.57 22.52 39.50
LLaVA-OV-0.5B [4] 9 29.48 30.14 49.22 42.72 18.04 14.92 31.48 25.67 28.98 30.10 15.89 24.43 21.75 1.50 33.42 50.88 50.00 32.00 27.75 26.04 30.91 34.01 24.50 24.65
LLaVA-OV-7B [4] 5 31.20 21.79 30.33 26.94 18.58 13.87 10.43 13.64 31.24 29.29 26.13 38.68 30.25 9.47 40.14 56.47 55.06 37.25 48.63 38.23 30.38 33.72 26.49 35.01
Qwen2-VL-2b [45] 11 24.60 19.43 38.03 40.63 18.84 14.09 7.81 7.07 17.82 11.14 27.55 26.21 25.25 31.20 28.22 54.12 49.11 21.75 25.27 12.47 23.92 27.62 24.83 14.85
Qwen2-VL-7b [45] 6 30.74 27.52 35.97 35.22 20.83 12.88 28.68 29.95 28.21 28.45 20.44 35.37 20.25 5.69 37.03 59.71 52.38 30.25 38.46 41.00 22.04 28.49 22.52 38.38
Qwen2.5-VL-7b [46] 2 33.07 28.75 31.33 33.66 21.99 14.97 42.88 37.73 23.83 23.64 22.97 33.33 28.75 6.83 40.27 58.24 51.49 44.75 50.00 32.13 33.87 32.85 27.15 31.93
LLaVA-v1.5-7b [48] 12 23.65 10.85 5.17 12.53 17.37 11.34 7.25 5.26 18.73 9.12 26.50 24.43 26.75 28.31 34.09 51.18 52.38 34.25 24.18 26.87 34.68 29.94 22.52 30.81
LLaVA-v1.6-7b [49] 13 13.21 8.53 12.14 0.00 20.35 0.27 10.76 0.41 24.27 0.00 4.79 6.62 7.75 0.00 20.18 51.76 7.74 6.25 32.14 6.37 39.52 10.47 21.52 5.88
@@ 4.2 Evaluation on SPAR-Bench
Table 4 shows that human performance consistently exceeds all models across difficulty levels, with a clear margin. While the strongest baseline (Qwen2.5-VL-72B) achieves competitive results, it still falls far short of human-level performance—particularly in tasks requiring spatial reasoning across views or precise geometric estimation.
@@
Fine-tuning InternVL2.5-8B on SPAR-mix significantly boosts performance across all difficulty levels, reaching 65.53 (low), 63.01 (mid), and 61.29 (high). The largest gains occur in depth estimation and multi-view reasoning—e.g., View Change Inference improves from 6.32 to 37.14—demonstrating the effectiveness of structured spatial data in aligning geometric cues across viewpoints. Notably, the fine-tuned model surpasses human performance in depth prediction, and even the strongest pre-finetuning model exceeds humans on distance estimation. This is understandable, as these metrics are MRA-based and favor numerically optimized models over human intuition.
@@
SQA3D [14] ScanQA [13]
Methods 3D EM@1 BLEU-4 CiDEr
3D-LLM [8] - 12.0 69.4
Chat-3D v2 [50] 54.7 14.0 87.6
LEO [15] 50.0 13.2 101.4
LL3DA [9] - 13.5 76.8
Scene-LLM [12] 54.2 12.0 80.0
LLaVA-3D [10] 55.6 14.5 91.7
Video-3D LLM [51] 58.6 16.2 102.1
SPAR-mix 58.1 15.3 90.7
@@
ScanRefer [52] 37.3 24.3
MVT [53] 40.8 33.3
ViL3DRel [54] 47.9 37.7
3D-LLM [8] 30.3 -
Chat-3D v2 [50] 35.9 30.4
Grounded 3D-LLM [55] 47.9 44.1
LLaVA-3D [10] 54.1 42.4
Video-3D LLM [51] 58.1 51.7
@@ 4.3 Evaluation on 3D-specific tasks
We use two benchmarks: ScanQA, which focuses on object attributes and spatial relationships within a scene, and SQA3D, a spatial benchmark where the task is to answer specific questions based on a given viewpoint description within the scene. This requires the model to demonstrate spatial imagination. We treat this as a multi-view task during training. As shown in table 6, despite not incorporating 3D information into the model, it still achieves the competitive scores when compared to the 3D-enhanced LLMs.
@@
We evaluate the 3D grounding capability of our model using the ScanRefer dataset. ScanRefer requires the model to locate objects in a scene based on textual descriptions. As shown in the table 6, the values in parentheses represent the performance without proposal refinement. We observed that the VLM’s depth predictions may sometimes be inaccurate when the distance is large due to the absence of 3D information. However, after refining, the results perform well.
These results underline the potential of leveraging a VLM for grounding, showcasing its ability to perform well without the need for 3D information, while also tackling a more complex generative task.
CV-Bench [34] SPAR-Bench

### v6

@@ 4.1 Evaluation on existing 2D spatial benchmarks
CV-Bench tong2024cambrian.
The Cambrian Vision-Centric Benchmark (CV-Bench) is a 2,638-sample dataset for evaluating 2D spatial reasoning. It repurposes vision datasets with natural language questions and uses visual prompts (e.g., bounding boxes) to reduce ambiguity and ensure reliable evaluation.
VSI-Bench yang2024thinking.
The Visual-Spatial Intelligence Benchmark (VSI-Bench) evaluates spatial reasoning from egocentric video data, containing 5,000+ QA pairs from 288 real-world videos sourced from ScanNet dai2017scannet, ScanNet++ yeshwanth2023scannet++, and ARKitScenes baruch2021arkitscenes. It covers diverse indoor environments and leverages object-level annotations for question generation. VSI-Bench includes eight tasks across three categories: configurational reasoning (object counting, relative position, route planning), measurement estimation (object and room size, absolute distance), and spatiotemporal reasoning (appearance order in videos). All questions undergo manual review to ensure accuracy.
To evaluate the effectiveness of our dataset, we compare models trained on SPAR-mix (ours with EMOVA-2M) against those trained solely on EMOVA-2M. Since our dataset focuses specifically on spatial perception and reasoning, we incorporate a subset of QA data from EMOVA to maintain performance on general-purpose vision-language tasks. Note that the Base VLM has already been pretrained on extensive data with optimized training strategies, making it an unfair baseline for assessing the incremental gain. Therefore, we consider Base VLM + EMOVA-2M as the proper baseline. These results validate the effectiveness of our dataset design and indicate that spatial supervision can be introduced without compromising general QA performance. As shown in Tab. 2, our model achieves state-of-the-art results on CV-Bench 3D and VSI-Bench, demonstrating significant improvements in spatial understanding. While there are minor drops on general benchmarks such as MMBench and TextVQA—mostly within 1%—this is expected, as our data emphasizes spatial reasoning rather than fine-grained vision-language alignment.
VSI-Bench yang2024thinking CV-Bench   tong2024cambrian BLINK fu2024blink 3DSRBench ma20243dsrbench Seed-Image li2024seed MME fu2023mme MMBench liu2024mmbench RealWorldQA miyanishi2021sim2realqa TextVQA textvqa
@@
GPT-4v achiam2023gpt - 64.3 73.8 51.14 - 71.6 1927 75.0 61.4 77.4
GPT-4o achiam2023gpt 34.0 - - 60.04 45.3 77.1 2310 83.4 75.4 -
Cambrain-8B tong2024cambrian - 72.3 72.0 - - 74.7 1547 75.9 64.2 77.8
LLaVA-OV-7B li2024llava 25.3 - - 48.2 44.1 - 1998 80.8 66.3 -
InternVL2-8B chen2024far 34.6 - - - - 75.4 2215 81.7 - 77.4
@@
Table 2: Comparison of model performance on various spatial understanding and 2D general benchmarks. The base VLM is InternVL2.5-8B, and we use dynamic patch of 3. Note: Our main comparison is between +EMOVA-2M and +SPAR-mix, as both build upon the same Base VLM and general QA data. For detailed discussion, see section˜4.1.
QA source Hypersim roberts2021hypersim SUNRGBD song2015sun nuScenes caesar2020nuscenes
@@
We also find that pretraining with our dataset enhances the model’s spatial understanding across diverse, non-homogeneous datasets. Although our dataset is primarily derived from indoor scene datasets such as ScanNet dai2017scannet, ScanNet++ yeshwanth2023scannet++, and Structured3D zheng2020structured3d, the pretrained model shows consistent improvements across all tasks constructed with datasets different from ours in CV-Bench 3D, as shown in Tab. 3.
Notably, our method achieves an 18-point gain on the outdoor nuScenes dataset, highlighting strong generalization beyond indoor scenes. This gain stems from spatial reasoning emphasized in our data—many CV-Bench 3D questions involve relative distance, similar to DistI-OO and ObjRel. Since these tasks rely on 3D geometric cues rather than appearance or domain-specific semantics, the learned spatial priors transfer effectively across diverse environments.
@@ Methods
GPT-4o achiam2023gpt 3 36.39 29.25 53.80 45.00 15.00 13.60 37.40 34.40 23.40 24.40 24.93 30 16 28.80 45.11 64 64 58 46 46 32 44 30 22
Claude-3.7-Sonnet claude 5 21.77 25.43 41.00 45.40 11.20 12.20 42.60 19.60 26.00 5.40 7.33 16 6 0.00 23.33 40 48 22 36 14 12 20 6 12
Qwen2-VL-72B wang2024qwen2 4 35.62 35.28 45.40 49.80 13.80 10.00 54.60 49.40 36.80 22.40 23.39 42 18 10.16 40.00 60 68 50 38 44 18 28 18 36
Qwen2.5-VL-72B bai2025qwen2 2 39.40 35.35 53.20 46.80 17.80 29.00 49.60 57.40 14.40 14.60 23.05 40 16 13.16 48.44 74 74 60 56 50 20 34 24 44
@@
GPT-4o achiam2023gpt 3 38.11 36.88 51.22 44.69 21.21 19.33 41.40 44.90 36.34 35.96 26.49 27.74 25.25 19.99 43.80 65.00 64.88 44.75 50.82 43.21 29.84 32.56 27.81 35.29
GPT-4.1 achiam2023gpt 2 41.60 41.95 48.25 42.66 20.86 23.58 58.29 52.33 48.65 40.44 44.02 59.29 28.75 22.02 42.93 71.76 67.26 46.25 54.95 41.00 30.38 29.65 20.20 24.93
Doubao-1.5-vision-pro doubao 1 41.74 33.24 34.89 37.22 25.24 21.40 35.04 34.99 40.58 36.57 47.36 55.73 39.00 29.47 49.49 74.71 69.64 35.75 70.33 47.65 34.95 33.14 35.76 42.86
InternVL2-2B chen2024far 19 28.06 21.74 18.06 24.81 23.20 20.97 19.47 19.95 26.83 20.61 22.83 39.69 23.00 5.81 35.42 51.18 55.95 46.00 31.59 23.82 36.02 34.30 17.55 22.41
InternVL2-4B chen2024far 12 32.01 28.94 23.94 27.22 20.00 18.12 42.57 40.16 31.29 28.18 29.16 49.87 21.00 16.62 35.70 56.76 55.36 40.25 36.81 25.21 28.76 32.27 21.19 24.65
InternVL2-8B chen2024far 11 33.02 26.83 25.75 30.88 20.67 20.78 39.03 36.19 19.15 22.19 36.49 63.36 28.00 18.11 37.37 64.71 54.46 42.75 37.36 26.32 34.14 31.10 20.86 24.65
InternVL2.5-2B chen2024expanding 16 30.14 25.79 39.67 39.72 12.12 15.03 30.94 29.59 20.22 19.02 22.93 37.91 24.25 6.64 36.41 51.47 56.85 50.25 33.79 24.10 27.15 35.17 26.49 22.41
InternVL2.5-4B chen2024expanding 15 30.55 25.66 29.06 32.97 21.77 16.83 20.84 26.85 28.13 28.79 29.75 47.07 33.25 8.92 35.16 54.12 58.93 35.50 29.67 34.63 24.73 31.39 19.21 28.29
InternVL2.5-8B chen2024expanding 6 36.28 29.46 25.78 29.31 23.79 18.76 46.82 42.68 22.62 25.89 31.88 61.32 28.00 6.32 43.80 59.71 56.85 51.75 44.23 41.55 36.56 41.57 22.52 39.50
InternVL2.5-26B chen2024expanding 7 34.11 24.18 41.83 36.28 18.09 17.96 22.62 17.97 22.01 16.63 50.11 69.47 30.75 7.03 42.40 62.94 58.33 45.25 54.67 50.97 24.46 25.00 27.48 32.49
InternVL2.5-38B chen2024expanding 8 33.83 26.00 42.03 38.81 17.10 17.77 17.58 17.84 29.37 27.51 30.89 44.53 17.25 9.70 44.13 69.12 66.67 43.75 64.29 37.67 25.27 31.98 31.79 26.61
LLaVA-OV-0.5B li2024llava 17 29.48 30.14 49.22 42.72 18.04 14.92 31.48 25.67 28.98 30.10 15.89 24.43 21.75 1.50 33.42 50.88 50.00 32.00 27.75 26.04 30.91 34.01 24.50 24.65
LLaVA-OV-7B li2024llava 13 31.20 21.79 30.33 26.94 18.58 13.87 10.43 13.64 31.24 29.29 26.13 38.68 30.25 9.47 40.14 56.47 55.06 37.25 48.63 38.23 30.38 33.72 26.49 35.01
Qwen2-VL-2B wang2024qwen2 20 24.60 19.43 38.03 40.63 18.84 14.09 7.81 7.07 17.82 11.14 27.55 26.21 25.25 31.20 28.22 54.12 49.11 21.75 25.27 12.47 23.92 27.62 24.83 14.85
Qwen2-VL-7B wang2024qwen2 14 30.74 27.52 35.97 35.22 20.83 12.88 28.68 29.95 28.21 28.45 20.44 35.37 20.25 5.69 37.03 59.71 52.38 30.25 38.46 41.00 22.04 28.49 22.52 38.38
Qwen2.5-VL-7B bai2025qwen2 10 33.07 28.75 31.33 33.66 21.99 14.97 42.88 37.73 23.83 23.64 22.97 33.33 28.75 6.83 40.27 58.24 51.49 44.75 50.00 32.13 33.87 32.85 27.15 31.93
Qwen2.5-VL-32B bai2025qwen2 9 33.09 27.09 37.47 35.25 15.91 15.75 34.20 33.01 26.20 18.96 33.92 34.10 33.75 13.15 40.44 58.82 61.31 37.75 51.10 34.35 26.08 33.14 25.83 35.57
Qwen2.5-VL-72B bai2025qwen2 5 37.01 29.94 37.47 43.00 19.52 18.36 38.72 36.44 27.80 18.25 44.61 56.49 32.75 17.27 43.80 58.82 61.90 40.75 53.57 45.98 26.88 35.17 34.11 36.97
LLaVA-v1.5-7B liu2024improved 21 23.65 10.85 5.17 12.53 17.37 11.34 7.25 5.26 18.73 9.12 26.50 24.43 26.75 28.31 34.09 51.18 52.38 34.25 24.18 26.87 34.68 29.94 22.52 30.81
LLaVA-v1.5-13B liu2024improved 18 28.62 25.92 36.28 32.84 9.06 9.41 33.72 31.07 30.85 24.11 31.03 31.55 30.50 12.10 32.33 54.71 50.60 46.25 20.60 23.82 25.54 25.87 17.22 26.33
LLaVA-v1.6-7B liu2024llavanext 22 13.21 8.53 12.14 0.00 20.35 0.27 10.76 0.41 24.27 0.00 4.79 6.62 7.75 0.00 20.18 51.76 7.74 6.25 32.14 6.37 39.52 10.47 21.52 5.88
SpaceR-7B ouyang2025spacer 4 37.55 30.90 36.14 35.16 15.40 19.25 40.13 35.07 36.97 29.16 31.93 38.68 40.00 17.23 45.61 62.35 61.61 52.25 51.92 46.81 37.90 36.05 24.83 34.17
@@ 4.2 Evaluation on SPAR-Bench
Tab. 4 shows that human performance consistently exceeds all models across difficulty levels, with a clear margin. While the strongest baseline (Qwen2.5-VL-72B) achieves competitive results, it still falls far short of human-level performance—particularly in tasks requiring spatial reasoning across views or precise geometric estimation.
@@
Fine-tuning InternVL2.5-8B on SPAR-mix significantly boosts performance across all difficulty levels, reaching 65.53 (low), 63.01 (mid), and 61.29 (high). Gains are largest in depth and multi-view reasoning (e.g., View Change Inference: 6.32→37.14), indicating better cross-view geometric alignment. Notably, the fine-tuned model surpasses human performance in depth prediction, and even the strongest pre-finetuning model exceeds humans on distance estimation. This is understandable, as these metrics are MRA-based and favor numerically optimized models over human intuition.
@@
SQA3D ma2022sqa3d ScanQA azuma2022scanqa
Methods 3D EM BLEU-4 CiDEr EM
3D-LLM hong20233d - 12.0 69.4 20.4
Chat-3D v2 huang2023chat 54.7 14.0 87.6 -
LEO leo 50.0 13.2 101.4 21.5
LL3DA chen2024ll3da - 13.5 76.8 -
Scene-LLM fu2024scene 54.2 12.0 80.0 27.2
LLaVA-3D zhu2024llava 55.6 14.5 91.7 27.0
Video-3D LLM zheng2024video 58.6 16.2 102.1 30.1
SPAR-mix 58.1 15.3 90.7 27.7
@@
ScanRefer chen2020scanrefer 37.3 24.3
MVT huang2022multi 40.8 33.3
ViL3DRel ViL3DRel 47.9 37.7
3D-LLM hong20233d 30.3 -
Chat-3D v2 huang2023chat 35.9 30.4
Grounded 3D-LLM chen2024grounded 47.9 44.1
LLaVA-3D zhu2024llava 54.1 42.4
Video-3D LLM zheng2024video 58.1 51.7
@@ 4.3 Evaluation on 3D-specific tasks
We use two benchmarks: ScanQA, which focuses on object attributes and spatial relationships within a scene, and SQA3D, a spatial benchmark where the task is to answer specific questions based on a given viewpoint description within the scene. This requires the model to demonstrate spatial imagination. We treat this as a multi-view task during training. As shown in Tab. 6, despite not incorporating 3D information into the model, it still achieves the competitive scores when compared to the 3D-enhanced LLMs.
@@
We evaluate the 3D grounding capability of our model using the ScanRefer dataset. ScanRefer requires the model to locate objects in a scene based on textual descriptions. As shown in the Tab. 6, the values in parentheses represent the performance without proposal refinement. We observed that the VLM’s depth predictions may sometimes be inaccurate when the distance is large due to the absence of 3D information. However, after refining, the results perform well. These results underline the potential of leveraging a VLM for grounding, showcasing its ability to perform well without the need for 3D information, while also tackling a more complex generative task.
Model APE Mean $\downarrow$ APE P50 $\downarrow$ APE P90 $\downarrow$
Base VLM 2.22 1.86 4.12
SPAR-mix 1.05 0.64 2.45
Table 7: BEV coordinate prediction on ScanNet++ val. SPAR-mix reduces spatial error.
4.4 Probing Implicit 3D Representations
We probe whether the model learns implicit 3D structure beyond QA metrics. Following language-guided cognitive maps yang2024thinking, the model predicts bird’s-eye-view (BEV) 2D coordinates for multiple objects from multi-view RGB only, with no 3D supervision at inference. Each instance uses an observer-centric frame anchored at the first view: origin at the camera center projected to the ground plane, the y-axis along the viewing direction, and the x-axis to the observer’s right. The prompt states these rules and asks for meter-level coordinates in a fixed format.
APE Mean $\downarrow$ APE P90 $\downarrow$
Distance (m) #Objs Base VLM SPAR-mix Base VLM SPAR-mix
$[0,1)$ 990 0.74 0.70 1.03 1.61
$[1,2)$ 2635 1.40 0.73 1.88 1.77
$[2,3)$ 1619 2.39 1.04 2.86 2.25
$[3,5)$ 1225 3.66 1.46 4.50 3.26
$[5,7)$ 346 5.62 2.41 6.55 4.64
$[7,10)$ 62 7.62 4.43 8.42 7.38
$10,\infty)$ 5 12.71 11.57 16.46 15.26
Table 8: BEV APE across distance bins. SPAR-mix shows larger gains in medium-to-far ranges.
We build a validation set of 1,000 ScanNet++ samples, each with three images and 3–7 objects with known 3D positions. We evaluate object-level Average Position Error (APE) and compare our SPAR-mix model with a 2D-only base VLM (InternVL2.5-8B).
SPAR-mix shows lower localization error than the base model. By binning objects by distance from the main view, SPAR-mix consistently outperforms the base model, with larger gains at medium–far ranges where depth and occlusion are harder (e.g., mean APE: 1.46 vs. 3.66 at 3–5 m; 2.41 vs. 5.62 at 5–7 m; 4.43 vs. 7.62 at 7–10 m).
These results suggest that training with SPAR-mix yields a scene-consistent internal layout that aggregates cues across views to infer geometry, even without explicit 3D input at inference.
CV-Bench tong2024cambrian SPAR-Bench